square root(6x+5)=Square root (53)
Please show how to solve for x

Answer by sravancv(7)

square root(6x+5)=Square root (53)
squaring on both sides
so
we get
6x+5= 53
now adding -5 on both sides
6x+5-5=53-5
6x=48
now dividing by 6 on both sides
so
x=8